# Variational inequalities and smooth-fit principle for singular stochastic control problems in Hilbert spaces

### v1

@@ Appendix B Some technical results
For $p\geq 2$ as in (2.12), for any $T>0$ and for some $M>0$ it holds:
@@
    $\mathsf{E}\big{[}\sup_{t\in[0,T]}\big{|}X^{n;x,I}_{t}\big{|}_{H}^{p}\big{]}\leq M\Big{(}1+|x|_{H}^{p}+\mathsf{E}\Big{[}\Big{|}\int_{0}^{T}\big{|}\hat{\vartheta}_{s}\big{|}_{H}\mathrm{d}|I|_{s}\Big{|}^{p}\Big{]}\Big{)}$;
@@
    $\lim_{n\uparrow\infty}\mathsf{E}\big{[}\sup_{t\in[0,T]}\big{|}X^{n;x,I}_{t}-X^{x,I}_{t}\big{|}_{H}^{p}]=0.$
@@
We start by proving (i). By (B-1), the Burkholder-Davis-Gundy inequality (see Theorem 1.111 in [27]), the fact that $|e^{t\mathcal{A}_{n}}|_{\mathcal{L}(H)}\leq e^{-\frac{n\delta}{n+\delta}t}$ (cf. Equation (B-14) in [27]), and an estimate analogous to (2.10), we have, for some constant $M>0$ (possibly depending on $p$ and changing from line to line),
$\mathsf{E}\big{[}\sup_{t\in[0,T]}\big{|}X^{n;x,I}_{t}\big{|}_{H}^{p}\big{]}$$\leq$ $M\Big{(}|x|_{H}^{p}+\mathsf{E}\Big{[}\sup_{t\in[0,T]}\Big{|}\int_{0}^{t}e^{(t-s)\mathcal{A}_{n}}\sigma\mathrm{d}W_{s}\Big{|}_{H}^{p}\Big{]}$
$+\mathsf{E}\Big{[}\sup_{t\in[0,T]}\Big{|}\int_{0}^{t}e^{(t-s)\mathcal{A}_{n}}\hat{\vartheta}_{s}\mathrm{d}|I|_{s}\Big{|}_{H}^{p}\Big{]}\Big{)}$
$\leq M\Big{(}|x|_{H}^{p}+\mathsf{E}\Big{[}\int_{0}^{T}|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}\mathrm{d}s\Big{]}^{\frac{p}{2}}+\mathsf{E}\Big{[}\Big{|}\int_{0}^{T}\big{|}\hat{\vartheta}_{s}\big{|}_{H}\mathrm{d}|I|_{s}\Big{|}^{p}\Big{]}\Big{)}$
$\leq M\Big{(}1+|x|_{H}^{p}+\mathsf{E}\Big{[}\Big{|}\int_{0}^{T}\big{|}\hat{\vartheta}_{s}\big{|}_{H}\mathrm{d}|I|_{s}\Big{|}^{p}\Big{]}\Big{)},$
@@
$\mathsf{E}\Big{[}\sup_{t\in[0,T]}\big{|}X^{n;x,I}_{t}-X^{x,I}_{t}\big{|}_{H}^{p}\Big{]}$$\leq$ $M\Big{(}\sup_{t\in[0,T]}\big{|}\big{(}e^{t\mathcal{A}_{n}}-e^{t\mathcal{A}}\big{)}x\big{|}^{p}_{H}$
$+\mathsf{E}\Big{[}\sup_{t\in[0,T]}\Big{|}\int_{0}^{t}\Big{(}e^{(t-s)\mathcal{A}_{n}}-e^{(t-s)\mathcal{A}}\Big{)}\sigma\mathrm{d}W_{s}\Big{|}_{H}^{p}\Big{]}$
$+\mathsf{E}\Big{[}\sup_{t\in[0,T]}\Big{|}\int_{0}^{t}\Big{(}e^{(t-s)\mathcal{A}_{n}}-e^{(t-s)\mathcal{A}}\Big{)}\hat{\vartheta}_{s}\mathrm{d}|I|_{s}\Big{|}_{H}^{p}\Big{]}\Big{)}.$
The first two addends on the right-hand side of the latter display equation converge to zero as $n\uparrow\infty$ as in the proof of [27, Thm. 1.131]. In order to deal with the third addend, define
@@
which is such that $\psi_{n}(s)\to 0$ as $n\uparrow\infty$ $\mathsf{P}$-a.s. by [27, Prop. B.34]. Since now $|\psi_{n}(s)|\leq 2|\hat{\vartheta}_{s}|_{H}$, and, for any $T>0$, $\int_{0}^{T}|\hat{\vartheta}_{s}|_{H}\mathrm{d}|I|_{s}<\infty$ $\mathsf{P}$-a.s. by (2.3) and $\mathsf{E}[|\int_{0}^{T}|\hat{\vartheta}_{s}|_{H}\mathrm{d}|I|_{s}|^{p}]<\infty$ because $I\in\mathcal{I}$, the dominated convergence theorem gives
$\lim_{n\uparrow\infty}\mathsf{E}\Big{[}\sup_{t\in[0,T]}\Big{|}\int_{0}^{t}\Big{(}e^{(t-s)\mathcal{A}_{n}}-e^{(t-s)\mathcal{A}}\Big{)}\hat{\vartheta}_{s}\mathrm{d}|I|_{s}\Big{|}_{H}^{p}\Big{]}\leq\lim_{n\uparrow\infty}\mathsf{E}\Big{[}\sup_{t\in[0,T]}\Big{|}\int_{0}^{t}\psi_{n}(s)\mathrm{d}|I|_{s}\Big{|}_{H}^{p}\Big{]}=0$
@@
Let $p\geq 2$ as in (2.12). Then, for any $T>0$ and for some $M>0$ it holds:
$\mathsf{E}\big{[}\sup_{t\in[0,T]}\big{|}X^{x,I}_{t}\big{|}_{H}^{p}\big{]}\leq M\Big{(}1+|x|_{H}^{p}+\mathsf{E}\Big{[}\Big{|}\int_{0}^{T}\big{|}\hat{\vartheta}_{s}\big{|}_{H}\mathrm{d}|I|_{s}\Big{|}^{p}\Big{]}\Big{)}.$

### v2

@@ Appendix B Some technical results
For all $T>0$ there exists $M>0$ such that
@@
    $\mathsf{E}\big[\sup_{t\in[0,T]}\big|X^{n;x,I}_{t}\big|_{H}^{2}\big]\leq M\Big(1+|x|_{H}^{2}+\mathsf{E}\Big[\Big|\int_{0}^{T}\big|\hat{\vartheta}_{s}\big|_{H}\mathrm{d}|I|_{s}\Big|^{2}\Big]\Big)$;
@@
    $\lim_{n\uparrow\infty}\mathsf{E}\big[\sup_{t\in[0,T]}\big|X^{n;x,I}_{t}-X^{x,I}_{t}\big|_{H}^{2}]=0.$
@@
We start by proving (i). Estimating the Brownian integral in (B-1) using Theorem 1.112 in [28], and an estimate analogous to (2.11) for the integral involving the singular control, we have, for some constant $M>0$ (possibly depending on $T$ and changing from line to line),
$\mathsf{E}\big[\sup_{t\in[0,T]}\big|X^{n;x,I}_{t}\big|_{H}^{2}\big]$ $\leq$ $M\Big(|x|_{H}^{2}+\mathsf{E}\Big[\sup_{t\in[0,T]}\Big|\int_{0}^{t}e^{(t-s)\mathcal{A}_{n}}\sigma\mathrm{d}W_{s}\Big|_{H}^{2}\Big]$
$+\mathsf{E}\Big[\sup_{t\in[0,T]}\Big|\int_{0}^{t}e^{(t-s)\mathcal{A}_{n}}\hat{\vartheta}_{s}\mathrm{d}|I|_{s}\Big|_{H}^{2}\Big]\Big)$
$\leq M\Big(|x|_{H}^{2}+\mathsf{E}\Big[\int_{0}^{T}|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}\mathrm{d}s\Big]+\mathsf{E}\Big[\Big|\int_{0}^{T}\big|\hat{\vartheta}_{s}\big|_{H}\mathrm{d}|I|_{s}\Big|^{2}\Big]\Big)$
$\leq M\Big(1+|x|_{H}^{2}+\mathsf{E}\Big[\Big|\int_{0}^{T}\big|\hat{\vartheta}_{s}\big|_{H}\mathrm{d}|I|_{s}\Big|^{2}\Big]\Big),$
@@
$\mathsf{E}\Big[\sup_{t\in[0,T]}\big|X^{n;x,I}_{t}-X^{x,I}_{t}\big|_{H}^{2}\Big]$ $\leq$ $M\Big(\sup_{t\in[0,T]}\big|\big(e^{t\mathcal{A}_{n}}-e^{t\mathcal{A}}\big)x\big|^{2}_{H}$
$+\mathsf{E}\Big[\sup_{t\in[0,T]}\Big|\int_{0}^{t}\Big(e^{(t-s)\mathcal{A}_{n}}-e^{(t-s)\mathcal{A}}\Big)\sigma\mathrm{d}W_{s}\Big|_{H}^{2}\Big]$
$+\mathsf{E}\Big[\sup_{t\in[0,T]}\Big|\int_{0}^{t}\Big(e^{(t-s)\mathcal{A}_{n}}-e^{(t-s)\mathcal{A}}\Big)\hat{\vartheta}_{s}\mathrm{d}|I|_{s}\Big|_{H}^{2}\Big]\Big).$
The first two addends on the right-hand side of the latter display equation converge to zero as $n\uparrow\infty$ as in the proof of [28, Thm. 1.112]. In order to deal with the third addend, define
@@
which is such that $\psi_{n}(s)\to 0$ as $n\uparrow\infty$ $\mathsf{P}$-a.s. by [28, Prop. B.34]. Since now $|\psi_{n}(s)|\leq 2|\hat{\vartheta}_{s}|_{H}$, and, for any $T>0$, $\int_{0}^{T}|\hat{\vartheta}_{s}|_{H}\mathrm{d}|I|_{s}<\infty$ $\mathsf{P}$-a.s. by (2.3) and $\mathsf{E}[|\int_{0}^{T}|\hat{\vartheta}_{s}|_{H}\mathrm{d}|I|_{s}|^{2}]<\infty$ because $I\in\mathcal{I}$, the dominated convergence theorem gives
$\lim_{n\uparrow\infty}\mathsf{E}\Big[\sup_{t\in[0,T]}\Big|\int_{0}^{t}\Big(e^{(t-s)\mathcal{A}_{n}}-e^{(t-s)\mathcal{A}}\Big)\hat{\vartheta}_{s}\mathrm{d}|I|_{s}\Big|_{H}^{2}\Big]\leq\lim_{n\uparrow\infty}\mathsf{E}\Big[\sup_{t\in[0,T]}\Big|\int_{0}^{t}\psi_{n}(s)\mathrm{d}|I|_{s}\Big|_{H}^{2}\Big]=0$
@@
For any $T>0$, there exists $M>0$ such that
$\mathsf{E}\big[\sup_{t\in[0,T]}\big|X^{x,I}_{t}\big|_{H}^{2}\big]\leq M\Big(1+|x|_{H}^{2}+\mathsf{E}\Big[\Big|\int_{0}^{T}\big|\hat{\vartheta}_{s}\big|_{H}\mathrm{d}|I|_{s}\Big|^{2}\Big]\Big).$